OpenStudy (anonymous):

x^2+14x+20=832 for X

jimthompson5910 (jim_thompson5910):

Answers are \[\large x=-7+\sqrt{861} \ \ \text{or} \ \ x=-7-\sqrt{861}\] See attached image for how I got those answers.

OpenStudy (anonymous):

This does not factor. You may complete the square: \[x^2+14x+49=832-20+49\] \[(x+7)^2=861\] \[x=-7\pm \sqrt{861}\]